OpenStudy (anonymous):

Convert the equation 12x – 3y = 6 to slope-intercept form.

OpenStudy (ivettef365):

slope intercept form y = mx + b

OpenStudy (ivettef365):

just isolate the y

OpenStudy (anonymous):

You need to put 12x – 3y = 6 in the form of y = mx + b I will start it off. 12x – 3y = 6 – 3y = -12x + 6 \[ \frac{– 3y}{-3} = \frac{-12}{-3}x + \frac{6}{-3} \] can you finish it?